# Chak Tou Leong6, Jiale Kang, Jinjin Gu7

### v1

@@ 2.3 Models and Datasets
Our selection of base models is guided by three key criteria: (1) Following the standard RL paradigm, we select models that have undergone SFT as a cold start phase to ensure sufficient initial reasoning capabilities and reasoning format; (2) We employ models across different parameter scales to disentangle size-specific effects; and (3) We incorporate distinct model families to enhance the generalizability of our findings. Based on these principles, we utilize three reasoning models: DeepSeek-R1-Distill-Qwen-1.5B (deepseekai2025deepseekr1incentivizingreasoningcapability), Nemo-Openmath-1.5B (moshkov2025aimo2), and DeepSeek-R1-Distill-Qwen-7B (deepseekai2025deepseekr1incentivizingreasoningcapability).
@@ 3.3 Scaling on Stronger Models
To verify the generalizability of our findings, we scale our evaluation to the 7B parameter regime using DeepSeek-R1-Distill-Qwen-7B. As summarized in Table 6, the relative performance hierarchy observed in smaller models remains largely consistent at this larger scale. Both DoRA and LoRA+ achieve an overall average accuracy of $55.0\%$, outperforming the standard LoRA baseline ($54.8\%$). This consistent superiority suggests that the advantages of magnitude-direction decoupling (in DoRA) and optimized learning rate ratios (in LoRA+) are not mere artifacts of smaller model scales, but are intrinsic to the RLVR optimization landscape. Notably, DoRA maintains its lead across several challenging benchmarks, such as AMC ($83.1\%$) and AIME25 ($38.7\%$). These results reinforce our conclusion that for large-scale reasoning models, employing architecturally enhanced or optimization-aware adapters is more effective than relying on the standard LoRA formulation.

### v2

@@ 2.3 Models and Datasets
Our selection of base models is guided by three key criteria: (1) Following the standard RL paradigm, we select models that have undergone SFT as a cold start phase to ensure sufficient initial reasoning capabilities and reasoning format; (2) We employ models across different parameter scales to disentangle size-specific effects. Based on these principles, we utilize two reasoning models: DeepSeek-R1-Distill-Qwen-1.5B (deepseekai2025deepseekr1incentivizingreasoningcapability), and DeepSeek-R1-Distill-Qwen-7B (deepseekai2025deepseekr1incentivizingreasoningcapability).
@@ 3.3 Scaling on Stronger Models
To verify the generalizability of our findings, we scale our evaluation to the 7B parameter regime using DeepSeek-R1-Distill-Qwen-7B. As summarized in Table 6, the relative performance hierarchy observed in smaller models remains largely consistent at this larger scale. Both DoRA and LoRA+ achieve an overall average accuracy of $55.0\%$, outperforming the standard LoRA baseline ($54.8\%$). This consistent superiority suggests that the advantages of magnitude-direction decoupling (in DoRA) and optimized learning rate ratios (in LoRA+) are not mere artifacts of smaller model scales, but are intrinsic to the RLVR optimization landscape. Notably, DoRA maintains its lead across several challenging benchmarks, such as AMC ($83.1\%$) and AIME25 ($38.7\%$). These results reinforce our conclusion that for large-scale reasoning models, employing architecturally enhanced or optimization-aware awdapters is more effective than relying on the standard LoRA formulation.